According to the label on a bottle of concentrated hydrochloric acid, the contents are 36.0% HCl by mass and have a density of 1.18 g/mL.

What volume of it would you need to prepare 401 mL of 2.60 M HCl?
Answer #1

Given:

Mass% of HCl = 36.0%

Density = 1.18g/mL

The molarity of concentrated HCl:

The concentration on the bottle of HCl pertains to the concentration in percent by mass. If we have 1000 mL of a solution, we can determine the mass of solution as shown.

1000 mL×1.18 g1 mL = 1180 g1000 mL×1.18 g1 mL = 1180 g

Calculating for the mass of solute HCl present,

mass HCl = 0.36×1180 gmass HCl = 424.8 gmass HCl = 0.36×1180 gmass HCl = 424.8 g

Thus, we arrive with the molarity below.

[HCl] = 424.8 g36gmol1 L = 11.8 M

HCl volume you need to prepare 401 mL of 2.60 M HCl:

Using the dilution equation, we get

V1 = C2*V2/C1

= (2.60 M)(401mL)(11.8 M)

= 88.35mL

Thus, we take 88.35 mL of 11.8 M HCl and dilute the solution to 401mL.